OpenStudy (anonymous):

In an isosceles trapezoid, the longest base is 11", a leg is 5", and the height is 4". Find the length of the shorter base of the trapezoid.

OpenStudy (anonymous):

3,4,5 is a pythagorean triplet. Hence Short base = 11 -2(4)=11-8=3

OpenStudy (anonymous):

Whats the answer 2in. 3 in. 5 in. 6 in.